what is the bond order and formal charge in case   of SO42-    and PO43-

Grade:11

2 Answers

Mayank
4 years ago
bond order is given by (no. of bonds )/(no. of positions)
bond order of SO42-  is 6/4=1.5
bond order of PO43- is 5/4=1.25
 
formal charge of SO42- = net charge/no. of oxygen atoms=-2/4=-0.5
formal charge of PO43- = net charge / no. of oxygen atoms =-3/4
Khimraj
4 years ago
Bond order is equal to Total no of Bonds ( sigma as well as pi) in different resonating structures between any 2 atoms in a molecule/ Total no of resonating structures. In case of PO43- Bond order= 5/4= 1.25. Note: 5 is the total no of sigma and pi between P and O. 4 is the total no of resonating structures.